# Validation of ML-UQ calibration statistics using simulated reference values: a sensitivity analysis

### v1

@@ I Introduction
The use of probabilistic(Pernot2022c, ) or simulated(Rasmussen2023, ) reference values has been recently proposed to palliate the absence of reference value for a calibration statistic. A simulated reference value is estimated by applying the statistic to synthetic datasets containing the actual uncertainties and simulated errors generated from the uncertainties using a model generative distribution $D$.
@@
  • for validation, a confidence band can be estimated from simulated reference curves. For all error statistics, the width of the confidence band depends on $D$, leading to ambiguous diagnostics when $D$ is not constrained.
The present study considers these two points for other calibration statistics. More specifically, it aims to test how simulated reference values for calibration statistics are sensitive to the choice of a generative distribution, how this impacts the validation procedure, and whether the uncertainty on the simulated reference value is a good choice of metric for validation. The focus is limited to statistics linked to ML regression tasks: the correlation coefficient between absolute errors and uncertainties (CC), an average calibration statistic (the mean squared z-scores, ZMS), and two conditional calibration statistics (the ENCE, its ZMS-derived analog ZMSE).
@@ II.1 Probabilistic generative model
Note that $D$ should not be mistaken for the distribution of errors, which is a compound distribution between $D$ and the distribution of uncertainties. Let us assume that the errors are drawn from a distribution $D$ (Eq. 1) with a scale parameter $\sigma$ distributed according to a distribution $G$. The distribution of errors, H, is then a scale mixture distribution with probability density function
@@ II.2 Calibration statistics derived from the generative model
where the first term of the RHS has been identified as the mean squared uncertainty $<u_{E}^{2}>$. This expression can be compared to the standard expression for the variance
@@
where $RMSE$ is the root mean squared error $\sqrt{<E^{2}>}$ and $RMV$ is the root mean variance ($\sqrt{<u_{E}^{2}>}$). The RCE has been shown to be very sensitive to the upper tail of the uncertainty distribution and to become unreliable for a large portion of the studied ML-UQ datasets(Pernot2024_arXiv, ). It is therefore not considered directly here, but it is used in a bin-based statistic of conditional calibration, the Expected Normalized Calibration Error(Levi2022, )
@@
where $RCE_{i}$ is estimated over the data within bin $i$.
Pernot(Pernot2023a_arXiv, ) has shown that the ENCE reference value is not zero, and that it depends on the binning scheme. As demonstrated in Appendix B, the ENCE depends on both the dataset size and the number of bins and does not have a predefined reference value.
According to the binning variable, the ENCE might be used to test consistency (binning on $u_{E}$) or adaptivity (binning on input features).(Pernot2023d, ) Because of the limited range of each bin, the ENCE is not expected to suffer from the same problem as the RCE when the binning variable is the uncertainty.
@@
Note that Rasmussen et al.(Rasmussen2023, ) threat the NLL as if it had no predefined reference value and needed a simulated reference. This is not the case when the NLL is defined by Eq. 15, but it might be for other likelihood definitions. In the present case, for a given set of uncertainties, validation of the NLL is equivalent to the validation of the ZMS, with the additional constraint of a normal generative model.
@@
where $i$ runs over the $N$ bins and $ZMS_{i}$ is estimated with the data in bin $i$. The logarithm accounts for the fact that ZMS is a scale statistic (a ZMS of 2 entails a deviation of the same amplitude as a ZMS of 0.5). As for the ENCE, the ZMSE measures conditional calibration, i.e. consistency if the binning variable is the uncertainty or adaptivity if it is an input feature. Also, ZMSE depends on the dataset size and bin number (Appendix B), and doe not have a predefined reference value. As observed for the ZVE defined by Pernot(Pernot2023a_arXiv, ), one might expect the ZMSE to be more reliable than the ENCE for heavy-tailed uncertainty distributions.
@@ II.4 Validation
[…]
@@
$\zeta(\vartheta_{est},\vartheta_{ref},I)=\begin{cases}\frac{\vartheta_{est}-\vartheta_{ref}}{I_{BS}^{+}-\vartheta_{est}}&if\,(\vartheta_{est}-\vartheta_{ref})\leq 0\\ \frac{\vartheta_{est}-\vartheta_{ref}}{\vartheta_{est}-I_{BS}^{-}}&if\,(\vartheta_{est}-\vartheta_{ref})>0\end{cases}$ (20)
@@
$|\zeta|\leq 1$ (21)
which is strictly equivalent to the interval test (Eq. 19). In addition, $\zeta$ provides valuable information about the sign and amplitude of the agreement between the statistic and its reference value.
@@ II.4.3 Recommended validation workflow
A recommended validation workflow for a given statistic $\vartheta$ is shown in Fig. 1. Note that it is essential that $\vartheta$ should be first confirmed as being fit to the purpose of the study, notably for datasets with highly skewed and/or heavy-tailed uncertainty distributions(Pernot2024_arXiv, ).
@@
For statistics with a known reference value, one can apply directly the benchmark method (Eq. 22). In the absence of a predefined reference value for a statistic, one might generate a simulated one (Sect. II.4.2), but a weak point of this method is that one has to choose a generative distribution $D$ linking the errors to the uncertainties (Eq. 1).
If $D$ is well constrained, one may proceed to the estimation of the simulated reference value $\tilde{\vartheta}_{D,ref}$ and use it for validation (Eq. 23). In absence of constraints on $D,$ it is essential to estimate the sensitivity of $\tilde{\vartheta}_{D,ref}$ to $D$. For this, at least two alternative shapes for $D$ have to be considered, for instance $D=N(0,1)$ and $D=t_{s}(\nu)$, the unit-variance Student’s-t distribution is defined as
@@
One should use a value of $\nu$ small enough to provide a contrast with the normal distribution, but not too small, as it might generate problematic error sets with many outliers. A value of $\nu=6$ is found as a good compromise in the following. If the simulated reference values for these choices of $D$ differ more than their statistical uncertainty, the statistic should not be used for validation. Otherwise, one might proceed as in the case of a predefined $D$ (Eq. 23).
@@ III Applications
The validation approach presented above is applied to nine datasets extracted from the ML-UQ literature, with a focus on the shapes of the distributions of the errors, uncertainties and z-scores and their consequences for the validation process, followed by a thorough study of the sensitivity of the procedure to various parameters, such as the uncertainty distribution and the shape of the generative distribution.
@@ III.1 The datasets
Nine test sets, including errors and calibrated uncertainties, have been taken from the recent ML-UQ literature for the prediction of various physico-chemical properties by a diverse panel of ML methods. The selection rejected small datasets and those presenting duplicated properties. Note these uncertainties have been calibrated by a palette of methods with various levels of success. The datasets names, sizes and bibliographic references are gathered in Table 1, and the reader is referred to the original articles for details. In the following, a short notation is used, e.g. ’Set 7’ corresponds to the QM9_E dataset.
Set # Name Size ($M$) Reference
1 Diffusion_RF 2040 Palmer et al.(Palmer2022, )
2 Perovskite_RF 3834 Palmer et al.(Palmer2022, )
3 Diffusion_LR 2040 Palmer et al.(Palmer2022, )
4 Perovskite_LR 3836 Palmer et al.(Palmer2022, )
5 Diffusion_GPR_Bayesian 2040 Palmer et al.(Palmer2022, )
6 Perovskite_GPR_Bayesian 3818 Palmer et al.(Palmer2022, )
7 QM9_E 13885 Busk et al.(Busk2022, )
8 logP_10k_a_LS-GCN 5000 Rasmussen et al.(Rasmussen2023, )
9 logP_150k_LS-GCN 5000 Rasmussen et al.(Rasmussen2023, )
@@
Those same datasets were tested by Pernot(Pernot2024_arXiv, ) for average calibration by the ZMS statistic and the result is reported in the first column of Table 2. Only four of the nine sets passed the test for average calibration (Sets 1, 6, 7 and 9). In the same study, it was shown that using the RCE statistic (Eq. 10) for average calibration testing was problematic, and notably that this statistic was unreliable for the sets marked as “Failed RCE” in Table 2 (Sets 2, 3, 4, 6 and 7). This issue was traced back to the high sensitivity of the root mean variance (RMV) involved in the RCE expression to the upper tail of the uncertainty distribution.
Set # Calibrated Failed RCE $\beta_{GM}(u_{E})$ $k=\nu/2$ $\vartheta=\nu/2*\sigma^{2}$ $\nu=2k$ $\nu_{eff}$
1 $\checkmark$ 0.172 - - - -
2 $\checkmark$ 0.419 - - - -
3 $\checkmark$ 0.485 4. 52 7. 32E-01 9. 04 6. 8
4 $\checkmark$ 0.438 1. 57 2. 44E-01 3. 14 3. 3
5 0.113 21. 10 1. 82 42. 20 4. 0
6 $\checkmark$ $\checkmark$ - - - - -
7 $\checkmark$ $\checkmark$ 0.524 1. 81 1. 72E-04 3. 61 2. 2
8 0.231 23. 40 1. 91 46. 80 3. 9
9 $\checkmark$ 0.223 16. 90 3. 91E-01 33. 80 2. 9
Table 2: Tailedness of the uncertainty distributions: $\beta_{GM}$ skewness of the uncertainties and parameters of the fit of the squared uncertainties by $IG(k,\theta)$ distributions. $k$ is the shape parameter, assimilated to $\nu/2$ in the NIG model, and $\vartheta$ is the scale parameter, which absorbs the scale of $u_{E}^{2}$, noted $\sigma^{2}$. The missing values mark the sets with a bad fit by the $IG$ model or a non-representative $\beta_{GM}$. The values of the $\nu_{eff}$ parameters obtained from the fit of the errors distribution by the NIG model (Table 3) are reported for comparison with $\nu=2k$.
In order to have a clear view of the distributions involved in the generative model for the estimation of simulated references, the distributions of $u_{E,}$, $E$ and $Z$ for the nine datasets are characterized in the next sections.
III.1.1 Uncertainty distributions
Fig. 2 presents histograms of the logarithm of squared uncertainties for the nine datasets. $u_{E}^{2}$ has been fitted by an inverse gamma distribution $IG(k,\theta)$, to assess the pertinence of the NIG model defined above. The best fit curves are compared to the data histograms and show a reasonable agreement, except for Sets 1, 2 and 6. Uncertainties of Sets 1 appear to be bimodal, which cannot be modeled by the $IG$ distribution, while for Set 2, the $IG$ model is defeated by a long left tail of small uncertainties. Set 6 presents a combination of both features. The parameters of the fitted $IG(k,\theta$) distribution are reported in Table 2. One sees that if one excludes the problematic datasets, the shape parameter covers a large range, from 1.57 to 23.4, corresponding to skewness values varying from infinity to about 1 (skewness of the $IG$ distribution is not defined for $k\leq 3$).
Figure 2: Uncertainty distributions. The histograms of $log(u_{E}^{2})$ are compared to the fit of $u_{E}^{2}$ by an Inverse-Gamma distribution (blue line). The mean variance ($MV=<u_{E}^{2}>$) is marked by a red line and the median variance (MedV) by an orange line. The distance between $MV$ and $MedV$ characterizes the skewness of the distribution of squared uncertainties.
The mean variance and median variance were marked in Fig. 2. The distance between those values is an estimator of the skewness of the $u_{E}^{2}$ distribution. To be more quantitative and focus on the uncertainties instead of their square, skewness of the uncertainty distributions was quantified by $\beta_{GM}$, a skewness statistic robust to outliers based on the scaled difference between the mean and median (Groeneveld1984, ; Pernot2021, ). $\beta_{GM}$ varies between -1 and 1 and is null for symmetric distributions. One can see in Table 2 that all the datasets that challenge the RCE statistic have $\beta_{GM}$ values above 0.4, except for Set 6, for which skewness is not a pertinent descriptive statistic for such a bimodal distribution. One can see for the sets with a successful $IG$ fit that $\beta_{GM}$ is correctly anti-correlated to $k$.
III.1.2 Distributions of errors and z-scores
@@
The distributions of errors $E$ and z-scores $Z$ are characterized to assess their shapes and departure from normality. Table 3 reports the mean values, standard deviations, relative bias and effective number of degrees of freedom $\nu_{eff}$ resulting from the fit of the data by a scaled and shifted Student’s-t distribution. The smaller $\nu_{eff}$, the farther one is from a normal distribution. This fit accounts for a possible bias in the dataset, which is non-negligible (above 5 % of the standard deviation) only for Sets 9 for E, and Sets 8 and 9 for Z. Figs 3-4 compare the fits of the errors and z-scores by normal and Student’s-t distributions.
Set # $\mu_{E}$ $\sigma_{E}$ $b_{E}$ (%) $\nu_{eff,E}$ $\mu_{Z}$ $\sigma_{Z}$ $b_{Z}$ (%) $\nu_{eff,Z}$
1 0. 0033(81) 0. 3678(81) 1 3. 0 -0. 027(22) 0. 980(30) 3 6. 0
2 0. 0034(61) 0. 377(13) 1 1. 4 -0. 018(15) 0. 940(26) 2 3. 3
3 0. 008(11) 0. 4810(98) 2 6. 8 0. 002(23) 1. 058(18) 0 20. 1
4 0. 001(10) 0. 637(15) 0 3. 3 -0. 021(18) 1. 107(16) 2 9. 1
5 0. 0019(60) 0. 2713(61) 1 4. 0 0. 006(20) 0. 920(21) 1 3. 9
6 0. 0044(50) 0. 310(14) 1 1. 3 -0. 005(16) 0. 992(37) 1 1. 4
7 0. 00131(29) 0. 0341(48) 4 2. 2 0. 0174(84) 0. 9858(99) 2 4. 4
8 0. 0116(39) 0. 2786(52) 4 3. 9 0. 050(14) 0. 961(16) 5 3. 9
9 -0. 0424(22) 0. 1533(39) 28 2. 9 -0. 260(13) 0. 951(23) 27 3. 1
@@
Table 3: Summary statistics for the errors ($E$) and z-scores ($Z$): $\mu_{X}$ is the mean value of variable $X$, $\sigma_{X}$ its standard deviation, $b_{X}=100(\mu_{X}-0)/\sigma_{X}$ measures relative bias and $\nu_{eff,X}$ is the number of degrees of freedom resulting from the fit of the data by a scaled and shifted Student’s-t distribution.
@@
Figure 3: Error distributions (histograms) with normal (red line) and scaled and shifted Student’s-t (blue line) fits. For legibility, the histograms have been truncated to ±$\pm 3$ standard deviations, hiding a few outlying errors.
Figure 4: Same as Fig. 3 for z-scores.
The histograms and $\nu_{eff}$ values indicate that none of the the error distributions is close to normal, which is a common feature of error distributions(Pernot2020b, ; Pernot2021, ), and is consistent with the generative model and the uncertainty distributions of these datasets (Table 2). The largest value of $\nu_{eff,E}$ is 6.8 for Set 3, which is still a rather small value (a normal distribution corresponds to $\nu_{eff}=+\infty$, but values above 20-30 could be considered as close-to-normal). One can note also that some values of $\nu_{eff,E}$ are smaller than 2 (for Sets 2 and 6), corresponding to Student’s-t distributions with infinite variance.
@@
If one considers the z-scores, there is a notable trend for the values of $\nu_{eff}$ to be larger than their counterparts for the errors, except for a few sets (5, 6, 8) for which they are identical. The only set for which one gets notably closer to normality is Set 3, with $\nu_{eff,Z}=20.1$. Overall, one has still relatively small $\nu_{eff}$ values, rejecting the normality of z-scores for 8 out of nine datasets. Note that the normality of z-scores should not be used as a calibration criterion.

### v2

@@ I Introduction
The use of probabilistic(Pernot2022c, ) or simulated(Rasmussen2023, ) reference values has been recently proposed to palliate the absence of reference for a calibration statistic. A simulated reference value is estimated by applying the statistic to synthetic datasets containing the actual uncertainties and simulated errors generated from the uncertainties using a model generative distribution $D$.
@@
  • for validation, a confidence band can be estimated from simulated reference curves. For all error statistics, the width of the confidence band depends on $D$, leading to ambiguous validation diagnostics when $D$ is not constrained.
The present study considers these two points for other calibration statistics. More specifically, it aims to test (i) how simulated reference values for calibration statistics are sensitive to the choice of a generative distribution, (ii) how this impacts the validation procedure, and (iii) whether the uncertainty on the simulated reference value is a good choice of metric for validation. The focus is limited here to statistics linked to ML regression tasks: the correlation coefficient between absolute errors and uncertainties (CC), an average calibration statistic (the mean squared z-scores, ZMS), and two conditional calibration statistics (the ENCE, its ZMS-derived analog ZMSE).
@@ II.1 Probabilistic generative model
Note that $D$ should not be mistaken for the distribution of errors, which is a compound distribution between $D$ and the distribution of uncertainties. Let us assume that the errors are drawn from a distribution $D$ (Eq. 1) with a scale parameter $\sigma$, itself distributed according to a distribution $G$. The distribution of errors, H, is then a scale mixture distribution with probability density function
@@ II.2 Calibration statistics derived from the generative model
where the first RHS term of Eq. 4 has been identified as the mean squared uncertainty $<u_{E}^{2}>$. This expression can be compared to the standard expression for the variance
@@
where $RMSE$ is the root mean squared error $\sqrt{<E^{2}>}$ and $RMV$ is the root mean variance ($\sqrt{<u_{E}^{2}>}$). The RCE has been shown to be very sensitive to the presence of heavy tails in the uncertainty and error distributions and to be unreliable for a large portion of the studied ML-UQ datasets(Pernot2024_arXiv, ). It is therefore not considered directly here, but it is used in a bin-based statistic of conditional calibration, the Expected Normalized Calibration Error(Levi2022, )
@@
where $RCE_{i}$ is estimated over the data within bin $i$. According to the binning variable, the ENCE might be used to test consistency (binning on $u_{E}$) or adaptivity (binning on input features).(Pernot2023d, )
Pernot(Pernot2023a_arXiv, ) has shown that the ENCE reference value is not zero, and that it depends on the binning scheme. As demonstrated in Appendix B, the ENCE depends on the bins size and does not have a predefined reference value.
In case of heavy-tailed uncertainty distributions, the ENCE is not expected to suffer from the same problem as the RCE when the binning variable is the uncertainty(Pernot2024_arXiv, ). However, it shares the same sensitivity in the case of heavy-tailed error distributions or when the binning variable is not the uncertainty.
@@
Note that Rasmussen et al.(Rasmussen2023, ) treat the NLL as if it had no predefined reference value and needed a simulated reference. This is not the case when the NLL is defined by Eq. 15, but it might be for other likelihood definitions. In the present case, for a given set of uncertainties, validation of the NLL is equivalent to the validation of the ZMS, with the additional constraint of a normal generative model.
@@
where $i$ runs over the $N$ bins and $ZMS_{i}$ is estimated with the data in bin $i$. The logarithm accounts for the fact that ZMS is a scale statistic (a ZMS of 2 entails a deviation of the same amplitude as a ZMS of 0.5). As for the ENCE, the ZMSE measures conditional calibration, i.e. consistency if the binning variable is the uncertainty or adaptivity if it is an input feature. Also, the ZMSE depends on the bins size (Appendix B), and doe not have a predefined reference value. As observed for the ZVE defined by Pernot(Pernot2023a_arXiv, ), one might expect the ZMSE to be more reliable than the ENCE for heavy-tailed uncertainty distributions, but it shares its sensitivity to heavy-tailed error distributions(Pernot2024_arXiv, ).
@@ II.4 Validation
The validation protocol has been presented in Ref(Pernot2024_arXiv, ). The main points are summarized here. For a given dataset $(E,u_{E})$ and statistic $\vartheta$, one estimates the statistic over the dataset $\vartheta_{est}$ and a 95 % confidence interval $I_{BS}=\left[I_{BS}^{-},I_{BS}^{+}\right]$ by bootstrapping using the Bias Corrected Accelerated (BCa) method(DiCiccio1996, ). One can then test that the target value for the statistic, $\vartheta_{ref}$, lies within $I_{BS}$, i.e.
@@
$\zeta(\vartheta_{est},\vartheta_{ref},I_{BS})=\begin{cases}\frac{\vartheta_{est}-\vartheta_{ref}}{I_{BS}^{+}-\vartheta_{est}}&if\,(\vartheta_{est}-\vartheta_{ref})\leq 0\\ \frac{\vartheta_{est}-\vartheta_{ref}}{\vartheta_{est}-I_{BS}^{-}}&if\,(\vartheta_{est}-\vartheta_{ref})>0\end{cases}$ (20)
@@
$|\zeta(\vartheta_{est},\vartheta_{ref},I_{BS})|\leq 1$ (21)
which is equivalent to the interval test (Eq. 19). In addition, $\zeta$ provides valuable information about the sign and amplitude of the agreement between the statistic and its reference value.
@@ II.4.3 Recommended validation workflow
A recommended validation workflow for a given statistic $\vartheta$ is shown in Fig. 1. Note that it is essential that $\vartheta$ should be first confirmed as being fit to the purpose of the study, notably for datasets with highly skewed and/or heavy-tailed uncertainty and/or error distributions(Pernot2024_arXiv, ).
@@
For statistics with a known reference value, one can apply directly the benchmark method (Eq. 22). In the absence of a predefined reference value for a statistic, one might generate a simulated one (Sect. II.4.2), but a crucial point of this method is the choice of a generative distribution $D$ linking the errors to the uncertainties (Eq. 1).
If $D$ is well constrained, one may proceed to the estimation of the simulated reference value $\tilde{\vartheta}_{D,ref}$ and use it for validation (Eq. 23). In absence of constraints on $D,$ it is essential to estimate the sensitivity of $\tilde{\vartheta}_{D,ref}$ to $D$. For this, at least two alternative shapes for $D$ have to be considered, for instance $D=N(0,1)$ and $D=t_{s}(\nu)$, where the unit-variance Student’s-t distribution is defined as
@@
One should use a value of $\nu$ small enough to provide a contrast with the normal distribution, but not too small, as it might generate problematic error sets with very heavy tails and many outliers. From numerical experiments in a former study(Pernot2024_arXiv, ), a value of $\nu=6$ is found to be a good compromise. If the simulated reference values for these choices of $D$ differ more than their statistical uncertainty, the statistic is deemed to be over-sensitive to $D$ and should not be used for validation. Otherwise, one might proceed as in the case of a predefined $D$ (Eq. 23).
@@ III Applications
The validation approach presented above is applied to nine datasets extracted from the ML-UQ literature, with a focus on the shape of z-scores distributions and its consequences on the validation process. This is followed by a thorough study of the sensitivity of the validation procedure to various parameters, such as the uncertainty distribution and the shape of the generative distribution.
@@ III.1 The datasets
Nine test sets with a priori calibrated uncertainties have been extracted from the recent ML-UQ literature for the prediction of physico-chemical properties by a diverse panel of ML and UQ methods. These datasets were tested by Pernot(Pernot2024_arXiv, ) for average calibration by the RCE and ZMS statistics. The characterization of the squared uncertainty and squared error distributions by robust skewness and kurtosis statistics ($\beta_{GM}$ and $\kappa_{CS}$)(Crow1967, ; Groeneveld1984, ; Bonato2011, ; Pernot2021, ) was used to screen heavy-tailed distributions likely to cause reliability problems in the statistical validation of RCE and ZMS. It was found that $\beta_{GM}$ and $\kappa_{CS}$ were mostly redundant for these datasets, and only $\beta_{GM}$ is reported in Table 1. $\beta_{GM}$ is robust to outliers, varies between -1 and 1 and is null for symmetric distributions. For $u_{E}^{2}$ distributions, an upper safety limit of $0.6$ for $\beta_{GM}$ was established based on an Inverse Gamma distribution model, while for $E^{2}$ and $Z^{2}$ this limit is 0.8, based on a Fisher distribution model(Pernot2024_arXiv, ). Values exceeding these limits are noted in bold face in Table 1.
Set # Name Size ($M$) $\beta_{GM}(u_{E}^{2})$ $\beta_{GM}(E^{2})$ $\beta_{GM}(Z^{2})$
1 Diffusion_RF(Palmer2022, ) 2040 0.40 0.82 0.73
2 Perovskite_RF(Palmer2022, ) 3834 0.72 0.94 0.83
3 Diffusion_LR(Palmer2022, ) 2040 0.66 0.74 0.69
4 Perovskite_LR(Palmer2022, ) 3836 0.74 0.82 0.69
5 Diffusion_GPR_Bayesian(Palmer2022, ) 2040 0.19 0.78 0.79
6 Perovskite_GPR_Bayesian(Palmer2022, ) 3818 0.50 0.96 0.95
7 QM9_E(Busk2022, ) 13885 0.93 0.98 0.78
8 logP_10k_a_LS-GCN(Rasmussen2023, ) 5000 0.30 0.79 0.78
9 logP_150k_LS-GCN(Rasmussen2023, ) 5000 0.30 0.77 0.75
@@
In this previous study(Pernot2024_arXiv, ), the analysis of $u_{E}^{2}$ and $E^{2}$ distributions enabled also to show that applying the generative model (Eq. 1) led to reject the normality of $D$ for all the studied datasets. Typically, the errors are much more dispersed than what would be expected from a normal generative distribution. As this is a central point for the estimation of simulated reference values, the distribution of $Z$, which should reflect the empirical shape of $D$ for calibrated datasets is analyzed next.
III.1.1 Distributions of z-scores
@@
The fit of z-scores distributions by a scaled-and-shifted Student’s-t distribution is done by maximum likelihood estimation (Delignette2015, ). Table 2 reports the mean values, standard deviations, relative bias and the number of degrees of freedom $\nu_{Z}$ The smaller $\nu_{Z}$, the farther one is from a normal distribution. Note that this fit accounts for a possible bias in the dataset, which is non-negligible for Sets 8 and 9 (above 5 % of the standard deviation). The values of $\nu_{Z}$ reported here are globally consistent with the ones obtained previously(Pernot2024_arXiv, ) for the fit of the distributions of $Z^{2}$ by a Fisher-Snedecor $F(1,\nu_{Z^{2}})$ distribution, except for Set 9 for which the bias was not taken into account. Fig 2 shows the comparison of the best fits of z-scores by normal and scaled-and-shifted Student’s-t distributions.
Set # $\mu_{Z}$ $\sigma_{Z}$ $b_{Z}$ (%) $\nu_{Z}$ $\nu_{Z^{2}}$
1 -0. 027(22) 0. 980(30) 3 6. 0 7. 9
2 -0. 018(15) 0. 940(26) 2 3. 3 4. 9
3 0. 002(23) 1. 058(18) 0 20. 1 15. 1
4 -0. 021(18) 1. 107(16) 2 9. 1 8. 2
5 0. 006(20) 0. 920(21) 1 3. 9 2. 7
6 -0. 005(16) 0. 992(37) 1 1. 4 0. 9
7 0. 0174(84) 0. 9858(99) 2 4. 4 4. 0
8 0. 050(14) 0. 961(16) 5 3. 9 3. 7
9 -0. 260(13) 0. 951(23) 27 3. 1 20. 2
@@
Table 2: Summary statistics for the z-scores ($Z$): $\mu_{Z}$ is the mean value, $\sigma_{Z}$ the standard deviation, $b_{Z}=100(\mu_{Z}-0)/\sigma_{Z}$ measures relative bias and $\nu_{Z}$ is the number of degrees of freedom resulting from the fit of the data by a scaled and shifted Student’s-t distribution. The last column reports the shape parameter for the fit of $Z^{2}$ distributions by a $F(1,\nu_{Z^{2}})$ distribution(Pernot2024_arXiv, ).
@@
Figure 2: Z-scores distributions (histograms) with normal (red line) and scaled and shifted Student’s-t (blue line) fits. For legibility, the histograms have been truncated to ±$\pm 3$ standard deviations, hiding a few outlying values.
@@
The only sets for which one gets close to normality are Set 3, with $\nu_{Z}=20.1$, and to a lesser degree Set 4, with $\nu_{Z}=9.1$. Overall, one has rather small $\nu_{Z}$ values, rejecting unambiguously the normality of z-scores for 7 or 8 out of nine datasets. As a side effect, let us note that the normality of z-scores should not be used as a calibration criterion, unless $D$ is known to be normal.